can you solve
-x-y=-2
4x+5y=16
using subsitution method?

Answer by jim_thompson5910(35256)

Solve the first equation for y to get

-x-y=-2

-y = x-2

y = -x+2

Then plug this into the first equation

4x+5y=16

4x+5(-x+2)=16

4x-5x+10=16

Tell me what you get for x when you solve that equation above. Also, tell me what you get for y when you plug the solution for x into y = -x+2